### v1

# Density-aware Global-Local Attention Network for Point Cloud Segmentation

Abstract
3D point cloud segmentation has a wide range of applications in areas such as autonomous driving, augmented reality, virtual reality and digital twins. The point cloud data collected in real scenes often contain small objects and categories with small sample sizes, which are difficult to handle by existing networks. In this regard, we propose a point cloud segmentation network that fuses local attention based on density perception with global attention. The core idea is to increase the effective receptive field of each point while reducing the loss of information about small objects in dense areas. Specifically, we divide different sized windows for local areas with different densities to compute attention within the window. Furthermore, we consider each local area as an independent token for the global attention of the entire input. A category-response loss is also proposed to balance the processing of different categories and sizes of objects. In particular, we set up an additional fully connected layer in the middle of the network for prediction of the presence of object categories, and construct a binary cross-entropy loss to respond to the presence of categories in the scene. In experiments, our method achieves competitive results in semantic segmentation and part segmentation tasks on several publicly available datasets. Experiments on point cloud data obtained from complex real-world scenes filled with tiny objects also validate the strong segmentation capability of our method for small objects as well as small sample categories.
I Introduction
As one of the important research directions in the field of 3D computer vision, 3D point cloud semantic segmentation has a wide range of applications in autonomous driving, AR/VR, and digital twins, etc. With the proposal of Transformer [1] network, which is based on the attention mechanism, the point cloud segmentation network has been further improved. According to the processing scale of the attention mechanism, 3D Transformers can be classified into two main categories: global Transformers and local Transformers. The former directly performs attention computation on all input points for global feature extraction; the latter first divides the point cloud into multiple windows and performs attention computation within the windows to extract local features.
Global Transformer point cloud networks (PCT [2], 3CROSSNet [3], Point-BERT [4], etc.) tend to be applied only to object-level tasks and cannot effectively handle scene-level point cloud inputs due to the large amount of computation required by itself. Moreover, this type of method requires sampling when processing point clouds, which tends to ignore categories with fewer samples or small objects, and focuses more on the main part of the point cloud, thus tends to be more suitable for tasks such as classification.
Existing local Transformer networks (Point Transformer [5], Stratified Transformer [6], PPT-Net [7]) tend to divide the whole point cloud into windows of uniform size and perform the local attention. Due to the different distances of objects from the acquisition device in space and the difference in the amount of objects in different areas, the densities of different areas in point cloud tend to be unbalanced. And these networks mentioned above need to sample the point cloud before performing local attention, so some of the point in each local area would be randomly discarded due to the limited computational resources. As a result, the networks tend to learn poorly for some of the categories, especially the small object categories distributed in the dense area of the point cloud, thus affecting the recognition accuracy of the corresponding categories. In addition, the information interaction range of local attention networks is often limited within a single window, or the information interaction range is further improved by moving window operations such as Stratified Transformer [6], but the existing operations like shifting window are limited in terms of the enhancement of the effective receptive field.
Figure 1: Global attention carried out between local areas and local attention with different window sizes within local areas of different densities in proposed method.
Figure 2: Overview of proposed methods (taking outdoor scenarios as an example).
Aiming at the above problems, we propose a 3D point cloud segmentation method that fuses global attention with density-aware local attention. Unlike other methods that fix the local window size for attention of each network layer, the entire input point cloud is first adaptively divided into a controllable number of equal-sized local areas according to its true dimensional size. Based on the number of 3D points in each local area, as shown in Figure 1, the local density is calculated, and all local areas are divided into several discrete parts according to different densities. Then different local window sizes are determined for local areas belonging to different discrete parts in the local attention module, in order to density-aware processing as shown in Figure 2. At the same time, each local area is regarded as an independent whole, and an additional module is set up to calculate the global attention between different local areas. The global and local attention are computed in parallel, and the fusion between local and global attention is performed in the upsampling stage of the network, so that each 3D point can get the global-local feature update. In addition, a 3D category-response loss function is proposed, which allows the network to pay attention to categories with different sample sizes and to objects of different sizes more balanced among them.
The main contributions of this paper are:
  • A novel density-aware local attention module is proposed. In this module, we determine the window size based on the density distribution, the higher the density, the smaller the window and vice versa.
  • We present a global attention module based on performing inter-area feature updates. It is designed to model long-range dependencies where each local area is considered as an independent whole.
  • We introduce the category response loss. We use the category information embedded in the scene to constrain the network in the middle layer so that the network can respond to the presence of categories in advance.
  • Our method obtains excellent experimental results on both public datasets and self-collected datasets of real scenes. In experiments on the commonly used indoor dataset S3DIS [8], our method achieves excellent segmentation results in most categories.
II Related Work
II-A 3D Point Cloud Segmentation
The goal of the 3D point cloud segmentation task is to process the point cloud input to the network to obtain the category labels corresponding to each 3D point. Commonly used 3D point cloud segmentation methods can be classified into the following categories based on the principles of the network model and the data format of the input network: voxel-based methods [9, 10], projection-based methods [11, 12, 13, 14], MLP-based methods [15, 16, 17, 18, 19], pointwise convolution-based methods [20, 21], methods combining image processing [22, 23, 24], and methods based on the attention mechanism [2, 5, 25, 6, 26, 27, 28, 29, 30, 31, 7, 3, 4, 32, 33]. Voxel-based approaches discretize the 3D space into regular grids to enable 3D convolutions, but often struggle with trade-offs between resolution and efficiency. Projection-based methods transform 3D points into 2D representations, leveraging mature 2D CNN architectures for feature learning. MLP-based methods, such as PointNet and its successors, operate directly on raw point clouds, introducing innovative mechanisms to handle unordered data. Recent advances in pointwise convolution and attention mechanisms have further enhanced the ability to capture local and global geometric relationships.
II-B Network based on Transformer
Since Transformer [1] networks have a substantial increase in processing performance compared to CNN networks in their original field of Natural Language Processing (NLP), many articles in the vision field have attempted to use the Transformer-based network. In the field of 3D vision processing, the Transformer architecture is applied to similar tasks in 3D.
Figure 3: Visualisation diagram of local area density division of point cloud data (the visualisation of point cloud file density division in the indoor dataset S3DIS [8].
Most of the current point cloud transformer networks [5, 25, 6, 26] only consider local features without considering long-range dependencies, and are unable to better extract complete global context information, and including the above attention methods based on the transformer network architecture, although it expands the range of sensory field and information interaction to a certain extent, part of the long-range information is still not considered, and the window division also affects the performance. In real scenes such as oil field production, there are a large number of empty or weakly textured areas, and at the same time, there are small valves, railings, pipelines and other slender objects, and it is difficult for existing methods to perform point cloud segmentation better. Therefore, the proposed method is designed in terms of local window segmentation strategy, global feature update path, and loss function to further solve the problems existing in the current network.
III Methodology
The framework of the point cloud segmentation network fusing global attention and density-aware local attention proposed in this paper is shown in Figure 2. Firstly, in the point cloud local area division and density calculation module, the input point cloud is divided into local areas, the point cloud density is calculated for each local area and the corresponding local density value is assigned, then all the local areas are divided into multiple discrete parts according to the specific density, and the proposed Density-based Global-Local Attention module (DbGLA) is used for feature processing of point cloud data.
III-A Density-based Global-Local Attention
The proposed DbGLA module is mainly divided into two main parts, namely the local feature update module and the global feature update module: the former determines the window sizes inside the local areas based on their densities and computes the local attention between the internal windows within each local area; the latter treats each local area as an independent whole, and computes the global attention between different local areas of the whole scene.
III-A1 Local Feature Updates
Conventional Transformer-based 3D point cloud semantic segmentation networks [2, 5, 25, 6, 26], due to the large computational resources they require, need to first divide the input point cloud into grids of uniform size and sample the panels when there are too many points in them for local feature updating. However, the density of the point cloud data varies in different local areas, and setting a uniform grid size will cause the dense area of the overall point cloud to discard more 3D points compared to the sparse area when sampling, resulting in the loss of too many 3D points of smaller objects in the dense area. In this regard, the overall point cloud is first divided into multiple local areas of the same size. Then we set the number of local areas to be divided, and calculate the size of the local areas according to the number of areas and the real space occupied by the input point cloud. After that, we count the number of 3D points in each local area, and calculate the density of each local area according to the volume of the corresponding area. After calculating the density of each local area, all areas are sorted from sparse to dense density, divided into multiple discrete parts of density. We then assign a density value to the corresponding local area and all 3D points contained within it, with the sparsest area having the smallest density value and vice versa.
We set the same window size for local areas belonging to the same discrete part of the point cloud density, setting larger windows for areas with smaller density values and vice versa. As shown in Fig. 3 is the visualisation of the indoor scene point cloud data processed using the proposed local area density division module. The different colours represent the different density degree of the corresponding local areas of the point cloud, and it can be seen that the local areas containing special categories (e.g., columns, cabinets, etc., which are boxed by purple) are distinctly different from the density degree assigned to the neighbouring areas.
In order to increase the effective receptive field, we build cross-window information interaction. The local attention module designed is shown in Figure 4.
Figure 4: A schematic diagram of two connected local attention calculation modules.
After identifying and dividing the small windows within each local area, we introduce the density-aware local attention. For Query points within a specific window, the selection of Key points consists of two main parts, which are densely sampling points within the same window and sparsely sampling points within neighbouring windows as well as other windows with the same degree of density. In addition, the multi-head self-attention is independently used within each small window. Given that the number of heads for calculating the attention is $n_{h}$ and the dimension of each attention head is $n_{d}$, and assuming that the number of 3D points within the $i-$th small window is $n_{i}$. Then the feature dimensions of all the input belonging to the $i-$th small window can be obtained as $x\in\mathbb{R}^{(n_{d}\times n_{h})\times n_{i}}$. As in Equation 1 to Equation 4 is the computational procedure for calculating the local multi-head self-attention for the corresponding window:
$Q,K,V=Linear_{q,k,v}^{\mathbb{R}^{(n_{d}\times n_{h})\times n_{i}}\rightarrow\mathbb{R}^{n_{d}\times n_{h}\times n_{i}}}(x),$ (1)
$attn_{i,j,h}=softmax(Q_{i,h}\cdot K_{j,h}),$ (2)
$out_{i,h}=\sum_{j=1}^{n_{i}}(attn_{i,j,h}\times V_{j,h}),$ (3)
$o=Linear^{\mathbb{R}^{n_{d}\times n_{h}\times n_{i}}\rightarrow\mathbb{R}^{(n_{d}\times n_{h})\times n_{i}}}(out),$ (4)
where $Linear$ represents the linear layer operation and its superscripts represent the input and output dimensions of this linear layer. The $attn_{i,j,h}$ is the attention value computed by dot product of the vectors corresponding to the $i$th Query point and the $j$th Key point of the $h$th attention head in the window. The $out$ is then the aggregated features obtained by the computation. Finally, the features are projected back to the dimension consistent with the input features through an additional linear layer to obtain the final output feature $o$.
Moreover, we choose to adaptively compute and adjust the local area size, so that the total number of local areas for different inputs can be similar, which facilitates the subsequent computation of the global attention. Also the above operation avoids excessive computation due to too many local areas in the case of a large input point cloud.
III-A2 Global Feature Updates
For the global feature update process, we regard each local area as a whole that aggregates the position and feature information of the corresponding area to calculate the attention map between different local areas. Before calculating the global attention, the information for local area needs to be aggregated, and this aggregation process is mainly divided into two steps, the first step is the aggregation of 3D spatial location information, as shown in Equation 5; the second step is the aggregation of feature information of the local area, as shown in Equation 6:
$p_{i}^{{}^{\prime}}=\frac{\sum_{p_{j}\in\mathcal{P}_{i}}{p_{j}}}{N_{i}},$ (5)
$f_{i}^{{}^{\prime}}=\sum_{p_{j}\in\mathcal{P}_{i}}{(f_{p_{j}}\bigoplus e_{p_{j}})},~{}e_{p_{j}}=Linear_{enc}^{\mathbb{R}^{3}\rightarrow\mathbb{R}^{enc}}(p_{j}-p_{i}^{{}^{\prime}}),$ (6)
where $p_{i}^{{}^{\prime}}$ represents the spatial coordinates after local area aggregation. $f_{i}^{{}^{\prime}}$ is the aggregated feature of the corresponding area. $\bigoplus$ represents the vector concatenation operation. $p_{j}\in\mathcal{P}_{i}$ denotes that the 3D point $p_{j}$ belongs to the $i$th local area $P_{i}$, $N_{i}$ denotes the number of 3D points in the $i$th local area, $p_{j}$ and $f_{p_{j}}$ stand for the spatial coordinates of the corresponding 3D point and the input features of the corresponding point, $e_{p_{j}}$ represents the relative position encoding of the 3D point $p_{j}$ and the relative position encoding of the centre of gravity $p_{i}^{{}^{\prime}}$ of the local area to which it belongs.
In addition, after completing the global attention, the updated information needs to be assigned to 3D points, and again this process can be divided into two major parts, which are the assignment of spatial coordinates of the 3D points $p_{j}^{out}=p_{j}$, and the process of assigning global features of the 3D points as shown by Equation 7:
$f_{p_{j}}^{out}=Linear^{\mathbb{R}^{fea+enc}\rightarrow\mathbb{R}^{fea}}(f_{i}^{{}^{\prime}}),~{}p_{j}\in\mathcal{P}_{i},$ (7)
where $p_{j}^{out}$ and $f_{p_{j}}^{out}$ represent the spatial coordinates and feature information of the corresponding 3D points after the global feature update, respectively.
III-B Loss Function
When training the network, the loss function applied in this paper is mainly divided into two parts, which are the weighted cross-entropy loss function that adaptively determines the weights of the categories according to different datasets, and the category-response loss function oriented to 3D point cloud processing, and fuses the two kinds of loss functions according to the weights given by the design. When using the ordinary cross-entropy loss function for network model training, if the sample number of individual categories is too large, the loss function will make the network model pay more attention to the category, thus affecting the processing accuracy of the network model for the category with too few samples. In this regard, this paper adaptively calculates the number of 3D points in each category of samples in the training set for different experimental datasets, and calculates the weights of different categories in the calculation of the cross-entropy loss function, and the cross-entropy loss function after adding the weights is shown in Equation 8:
$\mathcal{L}_{WCE}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=0}^{C-1}w_{k}l_{ik}\log(p_{ik}),$ (8)
TABLE I: Quantitative comparison experiment and ablation experiment results of semantic segmentation of indoor point cloud dataset S3DIS [8] Area 5.
Method OA mAcc mIoU ceiling floor wall beam column window door table chair sofa bookcase board clutter
PointNet [15] - 49.0 41.1 88.8 97.3 69.8 0.1 3.9 46.3 10.8 59.0 52.6 5.9 40.3 26.4 33.2
KPConv [21] - 72.8 67.1 92.8 97.3 82.4 0.0 23.9 58.0 69.0 91.0 81.5 75.3 75.4 66.7 58.9
PCT [2] - 67.7 61.3 92.5 98.4 80.6 0.0 19.4 61.6 48.0 85.2 76.6 67.7 46.2 67.9 52.3
ST [6] 91.5 78.1 72.0 96.2 98.7 85.6 0.0 46.1 60.0 76.8 92.6 84.5 77.8 75.2 78.1 64.0
SpT [27] 89.5 77.3 68.9 92.6 97.7 83.5 0.2 42.0 60.6 67.1 88.8 81.0 73.2 86.0 63.1 60.0
SPoTr [28] 90.7 76.4 70.8 - - - - - - - - - - - - -
w/o DbGLA 91.6 76.8 71.0 95.2 98.6 87.0 0.0 46.6 61.1 74.7 89.9 85.3 78.8 68.7 73.2 63.9
w/ DbGLA 92.1 79.0 72.5 96.1 98.5 87.9 0.0 48.7 61.7 76.5 92.3 85.2 78.5 77.3 75.6 64.1
where $C$ represents the total number of categories. $N$ represents the total number of points. $w_{k}$ represents the adaptive weights of the $k$th category in calculating the cross-entropy loss, which is inversely proportional to the ratio of the number of samples of the corresponding category points to the total number of points extracted from the partially randomly selected scenarios multiple times. $l_{i}k$ represents the labelling of whether the $i-$th point belongs to the $k-$th category or not, when the $i-$th point belongs to the $k-$th category, the corresponding $l_{ik}$ value is $1$, otherwise it is 00. The $p_{ik}$ represents the probability that the network predicts the $k-$th category for the $i-$th point.
In addition to this, a novel 3D category-response loss function (CR-Loss) is proposed to enhance the network’s learning for semantic contexts as shown in Equation 9:
$\mathcal{L}_{3DCR}=-\frac{1}{N}\sum_{k=0}^{C-1}\sum_{i=1}^{N}(s_{ik}^{l}\log(s_{ik}^{p})+(1-s_{ik}^{l})\log(1-s_{ik}^{p})),$ (9)
where $s_{ik}^{l}$ represents the binary classification labels of the $i-$th point against the $k-$th category, if the $i-$th point belongs to the $k-$th category, the corresponding $s_{ik}^{l}$ value is $1$, otherwise it is 00. The $s_{ik}^{p}$ represents the binary prediction result for the $i-$th point against the $k-$th category.
Unlike pixel-level loss, CR-Loss contributes equally to both large and small targets, and tends to improve substantially in practice for segmentation processing performance for small targets. In the standard training process for 3D semantic segmentation, the network usually using cross-entropy loss computed between the prediction and ground truth for each point, and they may have difficulty understanding the context without global information. In order to standardise the training of the context encoding module, we introduce category response loss using the category information embedded in the scene during processing.
An additional fully-connected layer with a sigmoid activation function is constructed in addition to the coding layer of the proposed network in order to make individual predictions about the presence of object categories in the scene and learnt through binary cross-entropy loss, a setup that forces the network to understand the global semantic information at a very small additional computational cost. In subsequent experiments, we will see that the inclusion of 3D CR-Loss significantly improves the segmentation of small targets.
From combining the two loss functions, the final loss $\mathcal{L}_{all}$ are the sum of $\mathcal{L}_{WCE}$ and $\mathcal{L}_{3DCR}$ using a hyperparameter $\lambda$ as shown in Equation 10:
$\mathcal{L}_{all}=\lambda\mathcal{L}_{WCE}+(1-\lambda)\mathcal{L}_{3DCR}.$ (10)
IV Experiments
In order to verify the effectiveness of the proposed network approach, we choose to quantitatively compare it with existing representative methods on commonly used indoor 3D point cloud datasets, S3DIS [8], ScanNet [34], and qualitatively compares the segmentation results with the baseline network processing on an outdoor 3D point cloud dataset, Semantic3D [35]. The proposed network is also applied to the part segmentation task by using the ShapeNet [36] dataset for the corresponding experiments. In addition, we collect point cloud data from oil field production scenarios and perform segmentation experiments on these outdoor real field point clouds to verify the generality of the proposed method. The designed ablation experiments evaluate the effectiveness of the DbGLA module with the proposed CR-Loss.
IV-A Experimental settings
IV-A1 Comparison Methods
In this paper, the proposed methods are compared with existing representative methods applied to the task of semantic segmentation of 3D point clouds, respectively. Specifically, we choose MLP-based methods (PointNet [15], PointNet++ [16], RandLA-Net [18]), point convolution methods (KPConv [21]), and other Transformer network-based methods for attention mechanisms, which include the latest state-of-the-art methods that have been published in recent years: Point Cloud Transformer [2], Point Transformer [5], Stratified Transformer [6], Superpoint Transformer [27], SPoTr [28], SuperCluster [29], Point Transformer V2 [26] and OctFormer [37].
IV-A2 Datasets
In addition to the public dataset, we also conduct experiments in a real-scene dataset, which is the self-acquisition point cloud data collected from the real oil field scene by the air-ground fusion acquisition equipment consisting of 16-line radar, camera, UAV, IMU, etc. The point cloud involves a real space covering about 30,000 square metres, and contains more open weakly textured areas in the point cloud scene, as well as fewer samples of elongated objects in the categories of pipelines, fences, poles, road surfaces, containers, scaffolding, pumping machines, etc.
IV-A3 Experimental Details
Following the practice of new point cloud segmentation models in recent years [27, 28, 29], we use Overall Accuracy (OA), mean Accuracy (mAcc), mean Intersection over Union (mIoU) for semantic segmentation and Class mIoU, Instance mIoU for part segmentation. The deployment platforms for the experiments are all PyTorch, the computing platform for the indoor scene experiments is a server equipped with two Titan RTX (24G*2 video memory), and the computing platform for the outdoor scene experiments is a host computer equipped with an RTX4090 (24G video memory). The optimiser in the training process is selected as Adam, and the initial learning rate is set to 0.006. For the hyperparameter settings involved in the article, the number of discrete parts of the densities is generally set to 5 by default in the indoor scenario, while the number of discrete parts is set to 7 in the outdoor scenario; and the hyperparameter used in the loss function to balance the two loss terms is set to 0.5 by default. In addition, the number of iterations of the proposed method is set to 100 by default for experimental training.
IV-B Indoor Scene Experiment
TABLE II: Quantitative comparison experiment results of semantic segmentation of indoor point cloud dataset S3DIS [8] (6-fold validation).
Method OA mAcc mIoU
PointNet [15] 78.6 66.2 47.6
PointNet++ [16] 81.0 67.1 54.5
KPConv [21] - 79.1 70.6
RandLA-Net [18] 88.0 82.0 70.0
PT [5] 90.2 81.9 73.5
SuperCluster [29] - - 75.3
Ours 91.7 85.0 75.8
Figure 5: Qualitative experiment and ablation experiment results of semantic segmentation of outdoor point cloud dataset Semantic3D [35].
Table I summarises the results for the indoor scene dataset S3DIS [8] Area 5 experimental scenarios. The experimental results are listed as overall and category-level accuracy metrics for the comparison methods [21, 15, 2, 6, 27, 28] and for the proposed method. Bolded data represent the best results for the corresponding metrics, while underlining represents the second best results for the metrics. In the overall indexes, they are better than other comparative methods and state-of-the-art methods, and the proposed method category intersection ratio indexes are also the best in six of the 13 categories, while the other seven categories have also reached the sub-optimal level. Table II summarises the results under the six-fold cross-validation experimental scenario for the indoor scene dataset S3DIS [8]. The overall accuracy, average accuracy, and average cross-validation ratio values of the comparison method [21, 15, 16, 18, 5, 29], as well as the method of this paper, are listed in the experimental results. In addition, Table III shows the experimental results on the validation set of the ScanNet [34] dataset.
Compared to the MLP-based as well as the point convolution methods, the proposed method, as an attention-based method, the 3D points input to the network can have a larger effective sensory field, which makes its segmentation accuracy significantly better than these comparative methods. Compared to other methods, our proposed DbGLA module determines different sizes of local window sizes to compute the local self-attention for different local areas within the point cloud densities, so that more dense parts of the point cloud can be allocated more windows to compute the local attention, resulting in a higher segmentation accuracy for the categories of windows, chairs, sofas, bookcases, etc., which have a smaller sample of objects compared to the ceiling, floor, and so on. Our method obtains contextual information at the level of the whole scene by computing the global attention between local areas to enhance the long range modelling capability of the network, giving the proposed backbone network a stronger feature extraction capability and thus a higher overall segmentation accuracy.
TABLE III: Quantitative comparison experiment results of semantic segmentation of indoor point cloud dataset ScanNet [34].
Method mIoU
PointNet++ [16] 53.5
KPConv [21] 69.2
ST [6] 74.3
PTv2 [26] 75.4
OctFormer [37] 75.7
Ours 77.5
IV-C Outdoor Scene Experiment
The results in Figure 5 represent the results of qualitative comparison experiments in employing two partial scenes from the Semantic3D [35] validation set of outdoor scene datasets. The left column of the figure shows the segmentation results of this paper’s method without incorporating the semantic coding loss function CR-Loss, the middle column shows the segmentation results after incorporating both CR-Loss as well as the DbGLA module, and the right column shows the data truth values of the corresponding scenes.
It can be seen that the proposed method is more effective in segmenting outdoor scenes. Compared with the method without adding the proposed 3D CR-Loss, the method of adding the 3D CR-Loss based on the existing DbGLA module can balance the network’s attention among different categories and sizes of objects. As a result, in the experimental results, as shown in the categories with fewer samples such as grass, hardscapes, trees, shrubs, etc., which are marked with red circle in Figure 5, the segmentation effect of our method is significantly improved in the small-sample categories.
Figure 6: Qualitative segmentation results of the proposed method on self-collected point cloud data of actual oil field scenes.
TABLE IV: Quantitative comparison experimental results of partial segmentation of 3D point cloud dataset ShapeNet-Part [36].
Method Cls. mIoU Ins. mIoU
PointNet [15] 80.4 83.7
PointNet++ [16] 81.9 85.1
KPConv [21] 85.0 86.2
PCT [2] - 86.4
PT [5] 83.7 86.6
ST [6] 85.1 86.6
Ours 85.2 87.0
In addition to experiments on public dataset, we collect point cloud data in real scenes for real scenario segmentation experiments. As shown in Figure 6, it is the visualisation result demonstration of the proposed method for segmentation processing on point cloud data of multiple self-produced oil field scenes after pre-training in the outdoor public dataset. Through the visual demonstration, it can be seen that the proposed method can better achieve more equal attention to the large sample categories and small sample categories, and large objects and small objects when processing point clouds. Especially for the long and thin objects such as poles, railings, pipes, etc., which are more distributed in the actual scene of the oil field, they can be identified and processed better.
IV-D Part Segmentation Experiment
Table IV summarises the experimental results for the 3D point cloud partial segmentation task dataset ShapeNet [36], where the comparison methods [21, 15, 16, 2, 5, 6], and the values of the category-averaged intersection and merger ratios versus the instance-averaged intersection and merger ratios of this paper’s method are listed respectively. From this comparison experiment, it can be seen that the proposed method outperforms the other compared methods in a variety of metrics in the partial segmentation task, again thanks to the newly proposed DbGLA attention module. For local areas with different densities, our method uses windows of different sizes in order to compute the local attention within the corresponding local area, and with the addition of the proposed 3D CR-Loss, it further balances the network’s attention to the different parts of each object, thus improving the segmentation accuracy in the partial segmentation task.
IV-E Ablation Experiment
[…]
V Conclusions
In this paper, we propose a point cloud segmentation method that fuses density-aware local attention with global attention to build a balanced perception of local detailes and scene context. Using the proposed Category Response Loss, our method achieves a better balance among different categories and sizes of objects, so that it can better discriminate the harder-to-recognise small objects. Quantitative as well as qualitative experimental results on multiple publicly available datasets demonstrate the effectiveness of the proposed attention mechanism and the novel category response loss. The follow-up work can focus on analysing the shortcomings and corresponding improvement methods on researching the existing Transformer-based 3D semantic segmentation network for indoor scenes to the 3D segmentation task of outdoor scenes.